# Fast Monte Carlo Tree Diffusion: 100× Speedup via Parallel Sparse Planning

### v2

@@ 3 The Planning Horizon Dilemma
Despite its strong performance on long-horizon and complex planning tasks, MCTD suffers from a fundamental computational inefficiency. This is because, it requires repeated partial denoising operations to generate feasible subplans. Worse yet, as the planning horizon increases, the search space grows exponentially, resulting in significantly higher computational cost.
@@ 4.1 Parallel planning
To address this issue, we introduce the Redundancy-Aware Selection (RAS) mechanism. Our key observation motivating this design is that the time spent on the Selection step, among the four stages of MCTD, is negligibly small: as shown in Table 1, it only takes 0.05% time of the MCTD steps. This allows us to remove the Selection (and the Backpropagation) step from the parallelization pipeline and instead implement it as a lightweight serial operation.
@@ 5 Related works
Diffusion models [26] have demonstrated significant success in long-horizon trajectory planning, particularly in sparse-reward settings, by learning to generate plan holistically rather than forward modeling [12, 2, 1, 34, 17]. Since diffusion denoising is computationally intensive, many works address this limitation through hierarchical planning [5, 9, 6]. Chen et al. [4] tackles the limitation of diffusion planners in generating detailed short-term actions by incorporating value learning policies [28] for low-level control. Chen et al. [2] introduces causal noise scheduling with semi-autoregressive planning to improve performance in causality-sensitive tasks. Recent work has explored leveraging increased computational budgets to improve plan quality [31, 33]. Both Yoon et al. [31] and Zhang et al. [33] propose tree search methods for the denoising process. Yoon et al. [31] combines auto-regressive denoising with guidance scale-based action control for complex long-horizon planning, while Zhang et al. [33] develops a tree search method with learnable energy functions for value estimation, applying it to reasoning tasks such as Sudoku. However, as inference-time scaling in diffusion denoising has only recently been explored, optimizing efficiency in this context remains relatively understudied.
MCTS [7] has achieved notable success across various decision-making domains, particularly when integrated with learned policies or value networks [24, 23]. It has also been applied to Large Language Models (LLMs) to enhance reasoning required task performances [29, 30, 32]. To address MCTS’s inherent inefficiency in sequential search, various enhancements have been proposed, including parallel search methods [18], novel node selection policies [16], and visited-node penalization in parallel searches [24].
@@ 6 Experiments
We evaluate Fast-MCTD using tasks from the Offline Goal-conditioned RL benchmark (OGBench) [21], aligning with the setup in MCTD [31]. They includes tasks such as point and ant maze navigation with long horizons, robot arm multi-cube manipulation, and visually and partially observable mazes. Metrics reported are mean success rate (%) and planning time (sec), averaged over 50 runs (5 tasks $\times$ 10 seeds), including mean ± standard deviation. Detailed configurations and control settings are available in Appendix A.
@@ 6.3 Visual planning
Table 4 reveals that Fast-MCTD offers a substantial efficiency improvement (25-60× faster than MCTD), even outperforming it on larger mazes. The abstract trajectory planning through Sparse MCTD (S-MCTD) reduces rollout horizon complexity and improves credit assignment in latent transitions [17] as shown in Figure 3. Consequently, S-MCTD surpasses both MCTD and P-MCTD performance on larger maps, and Fast-MCTD’s integration of both sparse and parallel strategies maximizes efficiency. Replanning further enhances MCTD but at significant computational cost, whereas Fast-MCTD achieves superior performance with relatively minimal overhead, comparable to Diffuser-Replan.
@@ 6.4 Ablation studies
Parallelism degree and interval size. Here, we study two key hyperparameters: the parallelism degree (number of concurrent rollouts) and the interval size (steps skipped between generated states). Figure 4 shows the performance of Fast-MCTD in different configurations. Increasing the parallelism degree initially reduces planning time through efficient parallelization, but high degrees (above 200) degrade success rates due to delayed tree updates and selection redundancy. For interval size, larger interval sizes significantly boost efficiency but risk performance loss if the abstractions become too coarse for accurate low-level control [31].
@@ 7 Discussion
While our Fast Monte Carlo Tree Diffusion framework significantly improves performance, certain trade-offs exist. P-MCTD requires high-performance computing resources for optimal parallelization, but even with modest parallel capabilities, substantial efficiency gains are achievable on standard hardware. Our approach introduces additional hyperparameters (parallelism degree and interval size), but our ablation studies demonstrate these parameters are robust across a wide range of values with clear guidelines for effective settings. The computational benefits of Fast-MCTD substantially outweigh the minor overhead of parameter tuning, enabling practical deployment of diffusion-based planning in time-sensitive applications that were previously infeasible.
@@ 8 Conclusion
We introduced Fast Monte Carlo Tree Diffusion (Fast-MCTD), a scalable diffusion-based planning algorithm that resolves the planning horizon dilemma inherent to MCTD. By combining parallelized tree search (P-MCTD) with sparse sub-trajectory abstraction (S-MCTD), Fast-MCTD achieves significant inference-time speedups—up to 100× faster than MCTD in some tasks, while maintaining or exceeding performance across challenging long-horizon tasks. We demonstrated its effectiveness in maze navigation, robotic manipulation, and visual planning benchmarks. These gains are enabled by algorithmic innovations such as search-aware parallel rollouts and coarse-to-fine diffusion over abstract plans. While Fast-MCTD alleviates core efficiency bottlenecks, further improvements may be possible by integrating adaptive sparsity or learning-based guidance selection. Our findings highlight that structured reasoning from test-time scalability and efficient inference process are not mutually exclusive, opening new avenues for fast, deliberative decision-making in high-dimensional domains.
@@ A.1 Computation resources
The experiments were conducted on a computing server equipped with 8 NVIDIA RTX 4090 GPUs, 512GB system memory, and a 96-thread CPU. Model training required approximately 3-6 hours per model, while inference for each experimental evaluation took approximately up to 5 minutes.
@@
Figure 5: The task illustrations: (a) long-horizon maze navigation, (b) multi-object robotic manipulation requiring compositional planning, and (c) visual maze planning from high-dimensional raw RGB observations.
@@ A.2 Environment details
Following prior work [31], we evaluated our methods (P-MCTD, S-MCTD and Fast-MCTD) on challenging scenarios from the goal-conditioned task benchmark, OGBench [21], including long-horizon maze navigation, robot arm manipulation, and visually observable maze environments as illustrated in Figure 4. All evaluation tasks were unseen during training, requiring the models to generalize their learned planning capabilities to novel scenarios at inference time.
@@ A.2.1 Long-horizon maze
To assess performance on extensive planning horizons, we tested our methods on point-mass and ant robot navigation tasks across three increasingly complex environments: medium, large, and giant mazes. Models were trained on the navigate dataset containing long trajectories but without specific task knowledge. At inference time, models must leverage their learned capabilities to solve previously unseen maze configurations, testing their generalization to complex long-horizon planning problems.
@@ A.2.2 Robot arm manipulation
The robot manipulation tasks require compositional planning to move multiple objects in specific sequences. For example, when tasked with stacking objects in a predetermined order, the model must generate a coherent plan that respects these sequential constraints. As demonstrated in prior work [31], holistic diffusion-based planners often struggle with such compositional tasks, particularly without explicit search or reward guidance. We evaluated Fast-MCTD in this challenging domain, focusing on efficiency improvements within a replanning framework while maintaining MCTD’s strong performance on compositional tasks.
@@ A.2.3 Visual maze
To evaluate MCTD in high-dimensional, partially observable settings, previous work [31] utilized the Visual Antmaze task from OGBench [21]. In this task, the agent must navigate from an initial observation to a target, both represented as 64×64 RGB images with 3 color channels. The images are encoded into an 8-dimensional latent space using a Variational Autoencoder (VAE) [13], providing a compact representation for efficient planning.
@@ A.3 Baselines
  • Diffuser [12]: We implemented the standard Diffuser with classifier-guided generation [8] as our foundational baseline [31], providing comparable guidance mechanisms across all tested methods.
  • Diffuser-Replan: To strengthen this baseline, we extended Diffuser with periodic replanning capabilities. This variant regenerates complete plans at fixed intervals, allowing it to adjust to intermediate state observations and potentially improve performance through iterative refinement.
  • Diffusion Forcing [2]: This method represents a variant of diffusion planning that incorporates causal noise scheduling with tokenized trajectories but without explicit search mechanisms. Its default implementation includes replanning capabilities, making it a strong comparative baseline alongside MCTD.
  • MCTD [31]: As our primary baseline, MCTD represents the state-of-the-art in inference-time scaling for diffusion planning. It combines causal noise scheduling with tree-based search, demonstrating exceptional performance across diverse planning tasks. However, our investigation reveals significant computational inefficiencies, particularly in scenarios where its planning capabilities are most valuable.
@@ A.4 Model hyperparameters
While maintaining consistency with the hyperparameter configurations used in prior work [31], but we describe the detailed hyperparameters to improve the reproducibility. We highlight the key parameters specific to our methods while complete implementation details will be available.
@@
A.5 The evaluation details
Similar to the hyperparameters, many details are consistent with the settings established in prior work [31], while describing key experimental settings below.
@@ A.5.1 Maze navigation with point-mass and ant robots
For pointmaze tasks, $[0,0.1,0.5,1,2]$ is applied and $[0,1,2,3,4,5]$ is set in antmaze evaluation.
@@
$500$ planning horizon is applied for medium and large sized mazes and $1000$ for giant sized mazes.
@@
For pointmaze tasks, the heuristic controller [12] is applied and the diffusion-based value-learning policy [28] is utilized for antmaze [4]. For value-learning policy low-level controller acts to the given subgoal which is re-assigned after achieving with $10$ steps interval.
@@ A.5.2 Robot arm cube manipulation
$[1,2,4]$ is applied for each object. MCTDs applies the object-wise guidance, so the size of guidance set is increased as the number of objects increases in the environment.
@@
$200$ for single cube tasks and $500$ for other tasks.
@@
The value-learning policy [28] is applied as done in antmaze tasks.
Redundant plans between object-wise plans
Following previous work [31], we applied redundant plans to resolve inconsistencies between object-wise plans, such as attempting to move too far from the current position or trying to grasp an object while the robot is already holding another object.
A.5.3 Visual pointmaze
Guidance set.
For visual maze tasks, a guidance scale set of $[0,0.1,0.5,1,2]$ is applied.
Planning horizon.
A fixed planning horizon of 500 steps is used for medium and large-sized mazes.
Low-level controller.
Inverse Dynamics (InvDyna) models were trained following the MCTD framework [31] using offline datasets encoded by a pretrained VAE. These models predict the action $\hat{a}_{t}$ required to transition from the current state $s_{t}$ to the next state $s_{t+1}$, given the partially observable nature of the task. To capture velocity-related information, the InvDyna model conditions its predictions on both the previous state $x_{t-1}$ and the current state $s_{t}$, formalized as:
@@ A.5.3 Visual pointmaze
Guidance in POMDP.
To enhance spatial awareness in partially observable environments, a Position Estimator, pretrained on offline datasets, predicts the agent’s location in the VAE’s latent space. This module provides a coarse but informative positional signal.
@@ B.1 Leaf parallelization ablations
As shown in Table 11, leaf parallelization yields mixed results across different environment scales. In medium and large mazes, it shows modest efficiency improvements (7.4s vs. 8.5s and 6.7s vs. 7.4s, respectively) with minimal impact on success rates. However, in the giant maze environment, leaf parallelization significantly increases planning time (17.0s vs. 12.8s), demonstrating reduced efficiency precisely where computational optimization is most critical.
@@ B.2 Distillation of Diffusion Planner Ablations
Diffusion model distillation represents a promising alternative approach for optimizing denoising processes by training student models to predict the results of multiple denoising steps performed by a teacher model [22]. We conducted a preliminary investigation of this technique as another potential direction for optimizing MCTD.
Following the progressive distillation approach [22], we trained a student model to predict two-step denoised outputs from the teacher model. However, our experiments revealed significant challenges in trajectory quality, as illustrated in Figure 6. Upon analysis, we identified that MCTD’s causal noise scheduling—which applies different noise levels to different subplans—introduces unique complications for distillation approaches. While standard diffusion distillation typically assumes uniform noise levels across the entire generated content, MCTD’s varying per-subplan noise levels create a substantially more complex learning task for the student model.
Our findings suggest that effective distillation for MCTD would require specialized techniques capable of handling these heterogeneous noise distributions. This represents an intriguing avenue for future research, but we leave this as one of the future works.
@@ Appendix C Algorithms
13:        $expansions\leftarrow[\,]$triangleright Parallel expansion phase
14:        for $node\in nodes\_to\_expand$ do
@@
24:        $simulations\leftarrow[\,]$triangleright Parallel simulation phase
25:        for $(node,child)\in new\_children$ do

### v3

@@ 3 The Planning Horizon Dilemma
Despite its strong performance on long-horizon and complex planning tasks, MCTD suffers from a fundamental computational inefficiency. This is because it requires repeated partial denoising operations to generate feasible subplans. Worse yet, as the planning horizon increases, the search space grows exponentially, resulting in significantly higher computational cost.
@@ 4.1 Parallel planning
To address this issue, we introduce the Redundancy-Aware Selection (RAS) mechanism. Our key observation motivating this design is that the time spent on the Selection step, among the four stages of MCTD, is negligibly small: as shown in Table 1, it constitutes only 0.05% time of the MCTD steps. This allows us to remove the Selection (and the Backpropagation) step from the parallelization pipeline and instead implement it as a lightweight serial operation.
@@ 5 Related works
Diffusion models [26] have demonstrated significant success in long-horizon trajectory planning, particularly in sparse-reward settings, by learning to generate plan holistically rather than forward modeling [12, 2, 1, 34, 17]. Given the computational intensity of the denoising process, many approaches have focused on improving efficiency through methods like hierarchical planning [5, 9, 6]. To address the challenge of generating detailed, short-term actions, Chen et al. [4] integrated value-learning policies [28] for low-level control. Chen et al. [2] introduces causal noise scheduling with semi-autoregressive planning to improve performance in causality-sensitive tasks. More recently, research has explored leveraging increased computational budgets at inference time to enhance plan quality [31, 33]. Both Yoon et al. [31] and Zhang et al. [33] employ tree search methods within the denoising process. Yoon et al. [31] combines auto-regressive denoising with guidance scale-based action control for complex long-horizon planning, while Zhang et al. [33] develops a tree search method with learnable energy functions for value estimation, applying it to reasoning tasks such as Sudoku. However, as inference-time scaling in diffusion denoising has only recently been explored, optimizing efficiency in this context remains relatively understudied.
Monte Carlo Tree Search (MCTS) [7] has a long history of success in decision-making domains, particularly when integrated with learned policies or value networks [24, 23]. It has also been applied to Large Language Models (LLMs) to enhance reasoning required task performances [29, 30, 32]. To address MCTS’s inherent inefficiency in sequential search, various enhancements have been proposed, including parallel search methods [18], novel node selection policies [16], and visited-node penalization in parallel searches [24].
@@ 6 Experiments
We evaluate Fast-MCTD using tasks from the Offline Goal-conditioned RL benchmark (OGBench) [21], aligning with the setup in MCTD [31]. These include tasks such as point and ant maze navigation with long horizons, robot arm multi-cube manipulation, and visually and partially observable mazes. Metrics reported are mean success rate (%) and planning time (sec), averaged over 50 runs (5 tasks $\times$ 10 seeds), including mean ± standard deviation. Detailed configurations and control settings are available in Appendix A.
@@ 6.3 Visual planning
As shown in Table 4, Fast-MCTD demonstrates a substantial efficiency improvement (25-60× faster than MCTD), even outperforming it on larger mazes. The abstract trajectory planning through Sparse MCTD (S-MCTD) reduces rollout horizon complexity and improves credit assignment in latent transitions [17] as shown in Figure 3. Consequently, S-MCTD surpasses both MCTD and P-MCTD on larger maps. By integrating both sparse and parallel search, Fast-MCTD maximizes these efficiency gains. Replanning further enhances MCTD but at significant computational cost, whereas Fast-MCTD achieves superior performance with relatively minimal overhead, comparable to Diffuser-Replan.
@@ 6.4 Ablation studies
Parallelism degree and interval size. We also study two key hyperparameters: the parallelism degree (number of concurrent rollouts) and the interval size (steps skipped between generated states). Figure 4 shows the performance of Fast-MCTD in different configurations. Increasing the parallelism degree initially reduces planning time through efficient parallelization, but high degrees (above 200) degrade success rates due to delayed tree updates and selection redundancy. For interval size, larger interval sizes significantly boost efficiency but risk performance loss if the abstractions become too coarse for accurate low-level control [31].
@@ 7 Discussion
While our Fast Monte Carlo Tree Diffusion framework significantly enhances performance, it introduces certain trade-offs. Optimal parallelization for P-MCTD benefits from high-performance computing, yet substantial efficiency gains are achievable even with modest parallel capabilities on standard hardware. Our approach also introduces new hyperparameters, namely the parallelism degree and interval size. However, our ablation studies confirm the robustness of these parameters and provide clear guidelines for their effective configuration. The computational advantages of Fast-MCTD substantially outweigh the minor overhead of parameter tuning, enabling the practical deployment of diffusion-based planning in time-sensitive applications where it was previously infeasible.
@@ 8 Conclusion
We introduced Fast Monte Carlo Tree Diffusion (Fast-MCTD), a scalable planning algorithm that leverages diffusion models to resolve the inherent planning horizon limitations of its predecessors. By combining a parallelized tree search (P-MCTD) with a sparse sub-trajectory abstraction (S-MCTD), our method achieves significant inference-time acceleration. In our experiments, Fast-MCTD performed up to 100$\times$ faster than the baseline MCTD while maintaining or exceeding performance on challenging long-horizon tasks, including maze navigation and robotic manipulation. These gains are driven by algorithmic innovations like search-aware parallel rollouts and a coarse-to-fine diffusion process over abstract plans. Although Fast-MCTD alleviates key efficiency bottlenecks, future work could explore further improvements by integrating adaptive sparsity or learned guidance selection. Ultimately, our findings demonstrate that test-time scalability and structured reasoning are not mutually exclusive, opening new avenues for developing fast, deliberative agents in high-dimensional domains.
@@ A.1 Computation resources
The experiments were conducted on a server equipped with 8 NVIDIA RTX 4090 GPUs, 512 GB of system memory, and a 96-thread CPU. Model training required approximately 3–6 hours per model, while inference for each evaluation took up to 5 minutes.
@@
Figure 5: Task Illustrations: (a) long-horizon maze navigation, (b) multi-object robotic manipulation requiring compositional planning, and (c) visual maze planning from raw RGB observations.
@@ A.2 Environment Details
Following prior work [31], we evaluate our methods (P-MCTD, S-MCTD, and Fast-MCTD) on challenging scenarios from the OGBench [21] goal-conditioned benchmark. These include long-horizon maze navigation, robot arm manipulation, and visual maze navigation, as illustrated in Figure 5. All evaluation tasks were unseen during training, requiring the models to generalize their planning capabilities to novel scenarios at inference time.
@@ A.2.1 Long-Horizon Maze
To assess performance over long planning horizons, we test our methods on point-mass and ant robot navigation in three increasingly complex environments: medium, large, and giant mazes. The models were trained on the navigate dataset, which contains long trajectories but lacks specific task information. At inference, the models must leverage their learned dynamics to solve these unseen maze configurations, thereby testing their generalization for complex, long-horizon problems.
@@ A.2.2 Robot Arm Manipulation
The robot manipulation tasks demand compositional planning to move multiple objects in a specific sequence. For instance, when tasked with stacking objects in a predetermined order, the model must generate a coherent plan that adheres to these sequential constraints. As shown in prior work [31], holistic diffusion-based planners often struggle with such compositional tasks without explicit search or reward guidance. We evaluated Fast-MCTD in this domain, focusing on improving efficiency within a replanning framework while preserving MCTD’s strong compositional planning performance.
@@ A.2.3 Visual Maze
Following Yoon et al. [31], we evaluate performance in high-dimensional, partially observable settings using the Visual Antmaze task from OGBench [21]. In this task, the agent navigates from an initial observation to a goal, where both are represented as $64\times 64$ RGB images. These images are encoded into an 8-dimensional latent space using a Variational Autoencoder (VAE) [13], providing a compact representation for efficient planning.
@@ A.3 Baselines
  • Diffuser [12]: We implemented the standard Diffuser with classifier-guided generation [8] as a baseline, ensuring comparable guidance mechanisms across all methods.
  • Diffuser-Replan: To strengthen the Diffuser baseline, we incorporated periodic replanning. This variant regenerates the entire plan at fixed intervals, allowing it to adapt to intermediate state observations.
  • Diffusion Forcing [2]: This approach utilizes causal noise scheduling with tokenized trajectories but omits an explicit search mechanism. Its default implementation includes replanning, making it a strong comparative baseline.
  • MCTD [31]: As our primary baseline, MCTD represents the state-of-the-art in inference-time scaling for diffusion planning. It integrates causal noise scheduling with a tree-based search. While effective, it exhibits significant computational inefficiencies in complex planning scenarios.
@@ A.4 Model Hyperparameters
We maintain consistency with the hyperparameter configurations from prior work [31] and describe the detailed hyperparameters here to enhance reproducibility. We highlight the key parameters specific to our methods; complete implementation details will be made available.
@@
A.5 Evaluation Details
Our experimental setup aligns with the hyperparameters and settings of prior work [31]. We outline the key configurations for each task below.
@@ A.5.1 Maze Navigation with Point-Mass and Ant Robots
For the point-mass maze tasks, we use a guidance scale set of $\{0,0.1,0.5,1,2\}$. For the antmaze evaluations, the set is $\{0,1,2,3,4,5\}$.
@@
We set the planning horizon to $500$ for medium and large mazes and $1000$ for giant mazes.
@@
For point-mass maze tasks, we employ the heuristic controller from Janner et al. [12]. For antmaze tasks, we utilize the diffusion-based value-learning policy from Chen et al. [4]. This value-learning policy acts as a low-level controller, navigating toward a given subgoal. A new subgoal is assigned every $10$ steps upon reaching the previous one.
@@ A.5.2 Robot Arm Cube Manipulation
We apply a guidance scale set of $\{1,2,4\}$ for each object. As MCTD employs object-wise guidance, the size of the total guidance set scales with the number of objects in the environment.
@@
The planning horizon is set to $200$ for single-cube tasks and $500$ for all other manipulation tasks.
@@
We use the same value-learning policy as in the antmaze tasks, following the approach in Wang et al. [28].
Redundant Plans
Consistent with Yoon et al. [31], we apply redundant plans to resolve inconsistencies between object-wise plans. This prevents actions such as attempting to move to a distant location or trying to grasp a new object while another is already held.
A.5.3 Visual Point-Mass Maze
Guidance Set
For the visual maze tasks, we use a guidance scale set of $\{0,0.1,0.5,1,2\}$.
Planning Horizon
A fixed planning horizon of $500$ steps is used for both medium and large mazes.
Low-Level Controller
Following the MCTD framework [31], we trained an Inverse Dynamics (InvDyna) model on offline datasets encoded by a pretrained VAE. To account for the partially observable nature of the task, the model predicts the action $\hat{a}_{t}$ required to transition from state $s_{t}$ to $s_{t+1}$ by conditioning on both the previous and current states, $(s_{t-1},s_{t})$, thereby capturing velocity information. This is formalized as:
@@ A.5.3 Visual Point-Mass Maze
Guidance in POMDP
To improve spatial reasoning in this partially observable environment, we use a Position Estimator pretrained on offline datasets. This module predicts the agent’s location within the VAE’s latent space, providing a coarse but effective positional signal for guidance.
@@ B.1 Leaf parallelization ablations
As shown in Table 11, leaf parallelization yields mixed results across different environment scales. Although leaf parallelization provides marginal efficiency gains in medium (8.5s vs. 7.4s) and large (7.4s vs. 6.7s) mazes, in the giant maze environment, leaf parallelization significantly increases planning time (17.0s vs. 12.8s), demonstrating reduced efficiency precisely where computational optimization is most critical.
@@ B.2 Distillation of Diffusion Planner Ablations
Diffusion model distillation is a prominent technique for accelerating the denoising process by training a student model to predict the output of multiple teacher steps [22]. We explored this as a potential optimization for MCTD.
Following the progressive distillation method [22], we trained a student model to predict the outcome of two teacher denoising steps. However, as shown in Figure 6, this yielded poor-quality trajectories. We attribute this to the causal noise scheduling inherent to MCTD, which applies different noise levels to different subplans. Standard diffusion distillation assumes a uniform noise level across the entire output, making MCTD’s heterogeneous, per-subplan noise levels a significantly more complex learning problem for the student model.
Our findings indicate that effectively distilling MCTD requires specialized techniques capable of handling these non-uniform noise distributions. We leave the development of such methods to future work.
@@ Appendix C Algorithms
13:        $expansions\leftarrow[\,]$
14:        for $node\in nodes\_to\_expand$ in parallel do $\triangleright$ Parallel expansion phase
@@
24:        $simulations\leftarrow[\,]$
25:        for $(node,child)\in new\_children$ in parallel do $\triangleright$ Parallel simulation phase